These are the changes identified on the scheme; they have not yet been tested, but welcome if anyone wants to do so.
First of all the general schema of the "Frog" Pixie with the changes:
The changes on the output filter:
The change of the crystal, near the mixer-oscillator NE602:
Changes in the input resonating filter: the ceramic resonator at 7023 khz is replaced by an inductor that resonates with the capacitor (22 pF) at 14060.
All must to be tested.